# The drastic impact of Eddington-limit induced mass ejections on massive star populations

### v1

@@ Abstract
Context. Massive stars emit copious amounts of radiation, profoundly affecting their environment in galaxies and contributing to the reionization of the Universe. However, their evolution and thus their ionizing feedback are still not fully understood. One of the largest gaps in current stellar evolution calculations is the lack of a model for the mass ejections that occur when the stars reach the Eddington limit, such as during an Luminous Blue Variable (LBV) phase.
@@
Methods. We employ the 1D stellar evolution code MESA, in which we implement a new mass-loss prescription that is acting when stellar models inflate too much when reaching the Eddington limit. We use lines of constant inflation factors in the HR diagram for a simple empirical calibration of the threshold value. Synthetic massive-star stellar populations using calculated grids of single-star models with this mass loss prescription are compared with the observed populations in the Large and Small Magellanic Clouds. In combination with already computed grids of binary evolution models, we investigate the impact of binarity on our predictions.
Results. Our single-star models reproduce key features of the observed stellar populations, namely (i) the absence of stars located beyond the Humphreys-Davidson limit, (ii) an upper limit of RSG luminosities, (iii) the faintest observed single WR stars, (iv) the absolute number of O-stars, WRs, and RSGs, (v) WO stars in low metallicity environments, and (vi) the positions of LBV stars in the HRD. We show that binarity still plays an important role in explaining the observed WR stars. However, a large fraction of the binary population can also be explained via self-stripping. Our binary population explains at the same time the 70% binary fraction of O-stars and the 40% binary fraction of WR stars. However, our synthetic population also has caveats, such as an overproduction of bright H-free WN stars.
Conclusions. Our results show that the effect of Eddington-limit induced mass ejections on the structure and evolution of massive stars can remove tension between predicted and observed massive star populations. A more fundamental treatment of these effects, in particular for hydrogen-poor stars, is needed to fully comprehend massive star evolution.
@@ 1 Introduction
Massive stars ($M_{\mathrm{ini}}\gtrsim 8\,M_{\odot}$), although relatively rare, are the powerful engines of the Universe. Through their intense ionizing radiation, strong stellar winds, LBV eruptions, and supernova explosions, they shape their surroundings, are the regulators of star formation, and sources of chemical enrichment within their host galaxies (e.g., mae1:83; dra1:03; hop1:14; ram1:18; cro1:19). Despite their significance, the evolution of massive stars, particularly at the high-mass end where their impact is the strongest, remains poorly understood. The predicted evolution of these stars is highly dependent on the assumptions about internal mixing processes and mass-loss rates applied during different evolutionary stages (e.g., she1:20; gil1:21; sab2:22; zap1:24).
@@
One of the major challenges facing current stellar evolution models is their tendency to overproduce extremely luminous red supergiant (RSG) stars (e.g., bro1:11; geo1:13; cho1:16; lim1:18; cos1:25) beyond the empirical Humphreys-Davidson (HD) limit ($\log L/L_{\odot}\gtrsim 5.5$), where no stars are observed (hum1:79; dav1:18). This issue has been subject to numerous studies, which explored the effects of mixing processes and different RSG mass-loss rates on the synthetic RSG populations across various metallicities (sch1:18; kle1:20; hig1:20; zap1:24; dec1:24). However, a solution to this problem remains elusive.
@@
LBVs are massive stars that have evolved into a region in the HRD where their outer atmosphere is barely gravitationally bound, leading to enhanced mass-loss rates with possibly episodic outbursts. During the LBV phase, a star can lose several solar masses, making it a critical component of massive stellar evolution that could potentially alleviate some of the tension of the overluminous RSG problem. However, our understanding of the LBV phase is limited. Its impact on the evolution of massive stars is unclear (smi1:26) and due to the absence of a consistent framework for modeling LBV eruptions, it is often ignored in detailed stellar evolution calculations. In recent years, progress has been made in understanding the physics, impact, and modeling of LBV mass ejections (gra1:21; muk1:24; che1:24). Nevertheless, a physical model that prevents stellar evolution simulations from surpassing the HD limit for a significant duration, especially at low metallicity ($Z$), is still lacking.
@@
A third major issue is the difficulty to explain the existence of the least luminous, apparently single Wolf-Rayet (WR) stars, particularly in metal-poor galaxies (she1:20). The strength of UV-radiation-driven stellar winds weakens (e.g., vin1:01; mok2:07; hai1:17; pau1:25), which leads to the expectation that the minimum luminosity of WR stars formed by single stars should increase with decreasing $Z$. Consequently, stellar evolution models predict that the faintest WR stars in low-metallicity environments are all post-interaction binary products (e.g., smi1:14; mas1:21; pau1:22). Contrarily, dedicated multi-epoch surveys of the least-luminous WR stars in low-$Z$ environments strongly suggest that these stars are currently single (foe1:03; sch1:24). To explain the presence of these faint single WR stars, single star stellar evolution calculations require enhanced mass-loss rates (e.g., yoo1:17; sab1:22). Yet, this approach is in contrast to empirical mass-loss measurements, which, over the past few decades, have exposed lower mass-loss rates for hot stars than those commonly assumed in standard mass-loss prescriptions (hai1:17; she1:19; pau1:25).
@@
The Large (LMC) and Small Magellanic Cloud (SMC) are two nearby galaxies with low metallicities (i.e., Fe-group elements) of $Z_{\mathrm{LMC}}=0.5\,Z_{\odot}$ and $Z_{\mathrm{SMC}}=0.14\,Z_{\odot}$, respectively (vin2:23). These galaxies have been subject to numerous observing campaigns, and their massive star content is mostly known. Particularly, their RSG populations (dav1:18; yan1:23) as well as the WR population, including their binary properties (neu1:18; bar1:01; bar2:01; foe1:03; hai1:14; hai1:15; she1:19), are mostly known. This makes these galaxies ideal testbeds to benchmark stellar evolution models.
@@ 2.1 Stellar evolution modeling
A detailed list of the input physics is provided in Appendix A, and our source files can be downloaded from Zenodo [^1] https://doi.org/10.5281/zenodo.17397155. Here, only the most relevant physics are mentioned. Our models incorporate semiconvective mixing with $\alpha_{\mathrm{sc}}=1$ (lan1:83; sch1:19), and step core overshooting with $f_{\mathrm{ov}}=0.345$ and $f_{\mathrm{ov},\,0}=0.01$ (bro1:11) during core hydrogen and helium burning. To account for stabilizing chemical gradients that may limit the growth of the overshooting regions, we include the Brunt-Vaisala frequency and set the stabilizing Brunt composition gradient to $B=0.1$. As a result, in most of our calculations, overshooting during core helium burning occurs primarily in the very early stages, before a steep chemical gradient between the convective core and the envelope forms.
@@ 2.2 Eddington-limit induced mass ejections
Figure 1: HRDs of the massive star population (symbols) in the LMC (left), and SMC (right). The different symbols mark the positions of observed RSG (small red plusses, dav1:18; yan1:23), YSG (crosses, neu1:10; neu1:12), LBVs (filled triangles with quiescent and outburst positions, hum1:16; kal1:18), OB stars (open circles and squares, eva1:04; tru1:04; tru1:05; hun1:08; bes1:11; bou1:13; urb1:17; cas1:18; sch2:18; ram1:18; ram1:19; duf1:19; bes1:20; bou1:21; ric1:22; pau1:23; ber1:24; gom1:25; alk1:25), partially stripped stars (white large plusses pau1:22; ric1:23; ram1:23; ram1:24), and WR stars (white stars for apparent single, gray stars for binaries cro1:02; hai1:14; hai1:15; tra1:15; she1:16; she1:19). The empirical HD limit (hum1:79) is indicated by the hatched regions. Solid black and gray lines show stellar evolution tracks of the Model no-Eject until core hydrogen depletion. For clarity only the tracks with initial masses $M_{\mathrm{ini}}=40\,M_{\odot}$, $63\,M_{\odot}$, $100\,M_{\odot}$, $158\,M_{\odot}$, and $251\,M_{\odot}$ are labeled and colored in black. Dotted black lines mark the H-ZAMS and He-ZAMS. Contours in the background highlight the degree of inflation covering $R/R_{\mathrm{core}}=1\,\text{--}\,3$. Inflations of $R/R_{\mathrm{core}}=1.6$, $2.2$, and $2.8$ for stars during hydrogen burning, and $R/R_{\mathrm{core}}=1.05$, $1.1$, and $1.15$ for the He-ZAMS models are marked by colored lines and numbers.
@@
A connection of the LBV phenomenon with the Eddington limit is evident since the pioneering work of lam1:88 and ulm1:88, who showed that the empirical HD-limit is essentially recovered by the Eddington limit in stellar model atmospheres. This connection is strengthened by 1D stellar structure and evolution models (ish1:99; gra1:12; koe1:15; san1:15; gra1:21), which found that stars approaching their Eddington limit locally in their envelopes inflate their radius to allow a sufficient radiative energy transport, a phenomenon typical for hydrostatic codes that cannot be recovered by 3D radiation-hydrodynamic calculations (jia1:15; jia1:18; pon1:21).
@@
In contrast to che1:24, who model eruptive mass loss in 1D massive star models based on the local super-Eddington luminosity, we choose an empirically calibrated estimate of the global instability of the inflated stellar envelopes. In the absence of systematic stability analyses of such envelopes, we assume here that a stronger inflation leads to a higher chance for instability. A possible instability mechanism is provided by strange mode pulsations (gau1:90) which have been shown to occur in supergiant stellar models near the Eddington limit (san2:13; lov1:14), and which have growth time scales that are close to the hydrodynamic time scale of the envelope. To approximate Eddington-limit induced mass ejections, we assume that a star can inflate to a certain degree without becoming unstable, but that there must be a threshold value after which a star will expel its envelope until it returns to a “stable” inflation (i.e., when inflation is again below the given threshold). In our models, we fully account for the effect of envelope inflation during core hydrogen burning. During core helium burning, we apply the local superadiabatic reduction method from MESA for numerical stability, which still allows for significant inflation.
@@
To calibrate the extent of inflation beyond which Eddington-limit induced mass ejections should be triggered, we calculated stellar evolution models up to core hydrogen depletion using the physical setup described in Sect. 2.1. In Fig. 1, we compare the model tracks to the observed stellar populations in the HR diagram, for the LMC and the SMC. Contours indicate the amount of inflation in our models across the HRD. Interestingly, the slope of lines with fixed inflation values is parallel to the HD limit, supporting our hypothesis that a certain amount of inflation in the stellar envelope might be stable. One can see that core hydrogen-burning stars exceed the HD limit when their inflation reaches $R/R_{\mathrm{core}}\approx 2$ (see also san1:17), which we adopt as the triggering criterion for Eddington-limit induced mass ejections for such stars.
@@
Based on these empirical limits, we define the maximum allowed inflation in our models as
@@
with a scaling factor $\alpha$ depending on the surface hydrogen abundance $X_{\mathrm{H}}$ as
@@
As shown by pet1:06, mass loss with rates above a critical value can reduce the size of the inflated envelopes. Therefore, to approximate the mass lost during an ejection episode, we adopt the critical mass-loss rate ($\dot{M}_{\mathrm{eject}}$) of pet1:06 needed to prevent a star from further inflating as
@@
where $M$ is the total mass of the star, $\rho_{\mathrm{min}}$ is the minimum density in the inflated layers. To avoid extreme jumps in the mass-loss rates, we increase the mass loss rate linearly between $R/R_{\mathrm{core}}=\alpha(X_{\mathrm{H}})\cdot 1.9$} and $R_{\mathrm{max}}/R_{\mathrm{core}}$. For illustrative purposes, we display in Fig. 12 stellar evolution tracks calculated with the Eddington-limit induced mass ejections and indicate their ejection episodes. One can see that the starting points of the ejection episode in the models overlap with the observed LBVs in the LMC and SMC.
We performed test calculations, where we varied the inflation threshold for the maximum inflation by $50\%$. While this slightly alters the number of models that exceed the HD limit and the maximum WR luminosity, it does not significantly affect the total time spent during the OB or WR phase of a specific model. This gives us confidence that our chosen approach allows us to study the impact of mass ejections, such as during an LBV phase, on stellar populations without being overly sensitive to the choice of the threshold parameters.
Also note, that the critical mass-loss rate and thus a star’s subsequent evolution is only weakly sensitive to the choice of stellar-wind prescription during hot-star evolutionary phases, since typical variations among commonly used OB- or WR-star mass-loss recipes primarily affect the integrated mass loss but are not high enough to significantly alter the local density structure and thus prevent the effect of envelope inflation, especially at low metallicity. However, when adopting different mass-loss prescriptions, the empirical limits used to define the maximum allowed degree of inflation may need to be recalibrated.

### v2

@@ Abstract
Context. Massive stars emit copious amounts of radiation, profoundly affecting their environment in galaxies and contributing to the reionization of the Universe. However, their evolution and thus their ionizing feedback are still not fully understood. One of the largest gaps in current stellar evolution calculations is the lack of a model for the mass ejections that occur when the stars reach the Eddington limit, such as during a Luminous Blue Variable (LBV) phase.
@@
Methods. We employed the 1D stellar evolution code MESA, in which we implement a new mass-loss prescription that becomes active when stellar models inflate too much when reaching the Eddington limit. We used lines of constant inflation factors in the Hertzsprung-Russell diagram (HRD) for a simple empirical calibration of the threshold value. We calculated synthetic massive-star stellar populations using grids of single-star models with this mass loss prescription compared them with the observed populations in the Large and Small Magellanic Clouds. Further, with already computed grids of binary evolution models, we investigated the impact of binarity on our predictions.
Results. Our single-star models reproduce key features of the observed stellar populations, namely, (i) the absence of stars located beyond the Humphreys-Davidson limit; (ii) an upper limit of red supergiant (RSG) luminosities; (iii) the faintest observed single Wolf-Rayet (WR) stars; (iv) the absolute number of O-stars, WRs, and RSGs; (v) WO stars in low metallicity environments; and (vi) the positions of LBV stars in the HRD. We show that binarity still plays an important role in explaining the observed WR stars. However, a large fraction of the binary population can also be explained via self-stripping. At the same time, our binary population explains the 70% binary fraction of O-stars and the 40% binary fraction of WR stars. However, our synthetic population also has caveats, such as an overproduction of bright H-free WN stars.
Conclusions. Our results show that the effect of the Eddington-limit induced mass ejections on the structure and evolution of massive stars can remove the tension between predicted and observed massive star populations. A more fundamental treatment of these effects, particularly for hydrogen-poor stars, is needed to fully comprehend massive star evolution.
@@ 1 Introduction
Massive stars ($M_{\mathrm{ini}}\gtrsim 8\,M_{\odot}$), although relatively rare, are the powerful engines of the Universe. Through their intense ionizing radiation, strong stellar winds, Luminous Blue Variable (LBV) eruptions, and supernova explosions, they shape their surroundings and are the regulators of star formation and sources of chemical enrichment within their host galaxies (e.g., mae1:83; dra1:03; hop1:14; ram1:18; cro1:19). Despite their significance, the evolution of massive stars, particularly at the high-mass end, where their impact is the strongest, remains poorly understood. The predicted evolution of these stars is highly dependent on the assumptions about internal mixing processes and mass-loss rates applied during different evolutionary stages (e.g., she1:20; gil1:21; sab2:22; zap1:24).
@@
One of the major challenges facing current stellar evolution models is their tendency to overproduce extremely luminous red supergiant (RSG) stars (e.g., bro1:11; geo1:13; cho1:16; lim1:18; cos1:25) beyond the empirical Humphreys-Davidson (HD) limit ($\log L/L_{\odot}\gtrsim 5.5$), where no stars are observed (hum1:79; dav1:18). This issue has been the subject of numerous studies exploring the effects of mixing processes and different RSG mass-loss rates on the synthetic RSG populations across various metallicities (sch1:18; kle1:20; hig1:20; zap1:24; dec1:24). However, a solution to this problem remains elusive.
@@
Luminous Blue Variables are massive stars that have evolved into a region in the Hertzsprung-Russell diagram (HRD) where their outer atmosphere is barely gravitationally bound, leading to enhanced mass-loss rates with possibly episodic outbursts. During the LBV phase, a star can lose several solar masses, making it a critical component of massive stellar evolution that could potentially alleviate some of the tension of the overluminous RSG problem. However, our understanding of the LBV phase is limited. Its impact on the evolution of massive stars is unclear (smi1:26), and due to the absence of a consistent framework for modeling LBV eruptions, it is often ignored in detailed stellar evolution calculations. Early pioneering modeling attempts tried to incorporate the effect of LBV eruptions in stellar evolution calculations, by increasing the mass-loss rates of stars massive enough to undergo an LBV phase as soon as they enter a specific evolutionary stage (such as the post-main sequence, van1:91) or to enhance the mass-loss in regions unstable against strange-mode pulsations (lan1:94). In recent years, progress has been made in understanding the physics, impact, and modeling of LBV mass ejections (gra1:21; muk1:24; che1:24). Nevertheless, a physical model that prevents stellar evolution simulations from surpassing the HD limit for a significant duration, especially at low metallicity ($Z$), is still lacking.
@@
A third major issue is the difficulty to explain the existence of the least luminous, apparently single WR stars, particularly in metal-poor galaxies (she1:20). The strength of UV radiation-driven stellar winds weakens (e.g., vin1:01; mok2:07; hai1:17; pau1:25), which leads to the expectation that the minimum luminosity of WR stars formed by single stars should increase with decreasing $Z$. Consequently, stellar evolution models predict that the faintest WR stars in low-metallicity environments are all post-interaction binary products (e.g., smi1:14; mas1:21; pau1:22). Contrarily, dedicated multi-epoch surveys of the least-luminous WR stars in low-$Z$ environments strongly suggest that these stars are currently single (foe1:03; sch1:24). To explain the presence of these faint single WR stars, single-star stellar evolution calculations require enhanced mass-loss rates (e.g., yoo1:17; sab1:22). Yet this approach is in contrast to empirical mass-loss measurements, which over the past few decades have exposed lower mass-loss rates for hot stars than those commonly assumed in standard mass-loss prescriptions (hai1:17; she1:19; pau1:25).
@@
The Large (LMC) and Small Magellanic Clouds (SMC) are two nearby galaxies with low metallicities (i.e., Fe-group elements) of $Z_{\mathrm{LMC}}=0.5\,Z_{\odot}$ and $Z_{\mathrm{SMC}}=0.14\,Z_{\odot}$, respectively (vin2:23). These galaxies have been subject to numerous observing campaigns, and their massive star content is mostly known. In particular, their RSG populations (dav1:18; yan1:23) as well as the WR population, including their binary properties (neu1:18; bar1:01; bar2:01; foe1:03; hai1:14; hai1:15; she1:19), are mostly known. This makes these galaxies ideal test beds to benchmark stellar evolution models.
@@ 2.1 Stellar evolution modeling
A detailed list of the input physics is provided in Appendix A, and our source files can be downloaded from Zenodo. [^1] https://doi.org/10.5281/zenodo.17397155 Here, only the most relevant physics are mentioned. Our models incorporate semiconvective mixing with $\alpha_{\mathrm{sc}}=1$ (lan1:83; sch1:19), and step core overshooting with $f_{\mathrm{ov}}=0.345$ and $f_{\mathrm{ov},\,0}=0.01$ (bro1:11) during core hydrogen and helium burning. To account for stabilizing chemical gradients that may limit the growth of the overshooting regions, we include the Brunt-Vaisala frequency and set the stabilizing Brunt composition gradient to $B=0.1$. As a result, in most of our calculations, overshooting during core helium burning occurs primarily in the very early stages, before a steep chemical gradient between the convective core and the envelope forms.
@@ 2.2 Eddington-limit induced mass ejections
Figure 1: Hertzsprung-Russell diagrams of the massive star population (symbols) in the LMC (left) and SMC (right). The different symbols mark the positions of observed RSG (small red plusses, dav1:18; yan1:23), YSG (crosses, neu1:10; neu1:12), LBVs (filled triangles with quiescent and outburst positions, hum1:16; kal1:18), OB stars (open circles and squares, eva1:04; tru1:04; tru1:05; hun1:08; bes1:11; bou1:13; urb1:17; cas1:18; sch2:18; ram1:18; ram1:19; duf1:19; bes1:20; bou1:21; ric1:22; pau1:23; ber1:24; gom1:25; alk1:25), partially stripped stars (white large plusses pau1:22; ric1:23; ram1:23; ram1:24), and WR stars (white stars for apparent single, gray stars for binaries cro1:02; hai1:14; hai1:15; tra1:15; she1:16; she1:19). The empirical HD limit (hum1:79) is indicated by the hatched regions. Solid black and gray lines show stellar evolution tracks of the Model no-Eject until core hydrogen depletion. For clarity only the tracks with initial masses $M_{\mathrm{ini}}=40\,M_{\odot}$, $63\,M_{\odot}$, $100\,M_{\odot}$, $158\,M_{\odot}$, and $251\,M_{\odot}$ are labeled and colored in black. Dotted black lines mark the H-ZAMS and He-ZAMS. Contours in the background highlight the degree of inflation covering $R/R_{\mathrm{core}}=1\,\text{--}\,3$. Inflations of $R/R_{\mathrm{core}}=1.6$, $2.2$, and $2.8$ for stars during hydrogen burning, and $R/R_{\mathrm{core}}=1.05$, $1.1$, and $1.15$ for the He-ZAMS models are marked by colored lines and numbers.
@@
A connection of the LBV phenomenon with the Eddington limit is evident since the pioneering work of lam1:88 and ulm1:88, who showed that the empirical HD-limit is essentially recovered by the Eddington limit in stellar model atmospheres. This connection is strengthened by 1D stellar structure and evolution models (ish1:99; gra1:12; koe1:15; san1:15; gra1:21), which found that stars approaching their Eddington limit locally in their envelopes inflate their radius to allow for a sufficient radiative energy transport, a phenomenon typical for hydrostatic codes that cannot be recovered by 3D radiation-hydrodynamic calculations (jia1:15; jia1:18; pon1:21).
@@
In contrast to che1:24, who model eruptive mass loss in 1D massive star models based on the local super-Eddington luminosity, we choose an empirically calibrated estimate of the global instability of the inflated stellar envelopes. In the absence of systematic stability analyses of such envelopes, we assume here that a stronger inflation leads to a higher chance for instability. A possible instability mechanism is provided by strange mode pulsations (gau1:90) which have been shown to occur in supergiant stellar models near the Eddington limit (kir1:93; san2:13; lov1:14), and which have growth time scales that are close to the hydrodynamic time scale of the envelope. To approximate Eddington-limit induced mass ejections, we assume that a star can inflate to a certain degree without becoming unstable, but that there must be a threshold value after which a star will expel its envelope until it returns to a “stable” inflation (i.e., when inflation is again below the given threshold). In our models, we fully account for the effect of envelope inflation during core hydrogen burning. During core helium burning, we apply the local superadiabatic reduction method from MESA for numerical stability, which still allows for significant inflation.
@@
To calibrate the extent of inflation beyond which Eddington-limit induced mass ejections should be triggered, we calculated stellar evolution models up to core hydrogen depletion using the physical setup described in Sect. 2.1. In Fig. 1, we compare the model tracks to the observed stellar populations in the HRD, for the LMC and the SMC. Contours indicate the amount of inflation in our models across the HRD. Interestingly, the slope of lines with fixed inflation values is parallel to the HD limit, supporting our hypothesis that a certain amount of inflation in the stellar envelope might be stable. One can see that core hydrogen-burning stars exceed the HD limit when their inflation reaches $R/R_{\mathrm{core}}\approx 2$ (see also san1:17), which we adopt as the triggering criterion for Eddington-limit induced mass ejections for such stars.
@@
Based on these empirical limits, we defined the maximum allowed inflation in our models as
@@
with a scaling factor, $\alpha$, depending on the surface hydrogen abundance, $X_{\mathrm{H}}$, as
@@
As shown by pet1:06, mass loss with rates above a critical value can reduce the size of the inflated envelopes. Therefore, to approximate the mass lost during an ejection episode, we adopted the critical mass-loss rate ($\dot{M}_{\mathrm{eject}}$) of pet1:06 needed to prevent a star from further inflating as
@@
where $M$ is the total mass of the star and $\rho_{\mathrm{min}}$ is the minimum density in the inflated layers. To avoid extreme jumps in the mass-loss rates, we increased the mass-loss rate linearly between $R/R_{\mathrm{core}}=\alpha(X_{\mathrm{H}})\cdot 1.9$} and $R_{\mathrm{max}}/R_{\mathrm{core}}$. For illustrative purposes, we display in Fig. 12 the stellar evolution tracks calculated with the Eddington-limit induced mass ejections and indicate their ejection episodes. One can see that the starting points of the ejection episode in the models overlap with the observed LBVs in the LMC and SMC.
We performed test calculations where we varied the inflation threshold for the maximum inflation by $50\%$. While this slightly alters the number of models that exceed the HD limit and the maximum WR luminosity, it does not significantly affect the total time spent during the OB or WR phase of a specific model. This gave us confidence that our chosen approach allows for the impact of mass ejections, such as during an LBV phase, on stellar populations to be studied without being overly sensitive to the choice of the threshold parameters.
We also note that the critical mass-loss rate, and thus a star’s subsequent evolution, is only weakly sensitive to the choice of stellar-wind prescription during hot-star evolutionary phases, since typical variations among commonly used OB- or WR-star mass-loss recipes primarily affect the integrated mass loss but are not high enough to significantly alter the local density structure and thus prevent the effect of envelope inflation, especially at low metallicity. However, when adopting different mass-loss prescriptions, the empirical limits used to define the maximum allowed degree of inflation may need to be recalibrated.